OpenStudy (anonymous):

Group Theory: any one pls guide me with a link to find the definition of these......... 1.(Zm,+) 2.(Zm,*)

OpenStudy (experimentx):

what is Zm?

OpenStudy (zarkon):

they prob mean \(Z_m=\{0,1,2,\ldots,m-1\}\)

OpenStudy (zarkon):

the set of integers mod \(m\)

OpenStudy (anonymous):

Yeah!!!! i need complete sentences to explain it... do u have any stuff???

OpenStudy (experimentx):

show that it satisfies the property to be called a group.

OpenStudy (anonymous):

Exactly!!! can i get that proof or its details... i mean any link???

OpenStudy (experimentx):

let 'a' and 'b' be two elements of Zm, show that it is closed in +, ie, \( (a+b) \mod b \in Z_m \)

OpenStudy (experimentx):

addition operation is associative, and find the identity element of Zm

OpenStudy (experimentx):

Edit:: \[ (a+b) \mod m \in Z_m \]

OpenStudy (experimentx):

this is always true becuse \( (a+b) \mod m = c < m, c \in Z\m \)

OpenStudy (anonymous):

What i have written below is correct right?? or should there be any change????? Definitions: Let Zm be the set of non negative integers less than m: {0,1, ., m−1} The operation +m is defined as a +m b = (a + b) mod m. This is addition modulo m. The operation ∙m is defined as a ∙m b = (a + b) mod m. This is multiplication modulo m. Using these operations is said to be doing arithmetic modulo m. Example: Find 7 +11 9 and 7 ·11 9. Solution: Using the definitions above: – 7 +11 9 = (7 + 9) mod11 = 16 mod11 = 5 – 7 ·11 9 = (7 ∙ 9) mod11 = 63 mod 11 = 8

OpenStudy (experimentx):

huh!! what exactly is the question?

OpenStudy (anonymous):

I wanna define it with examples!

OpenStudy (experimentx):

aren't those two different distinct groups?

OpenStudy (anonymous):

u mean the addition modulo and the other???

OpenStudy (experimentx):

yep!! just this one .. (Zm,+), this is a group.

OpenStudy (anonymous):

Thats it...what is confusing u??

OpenStudy (experimentx):

you are mixing two things ... doing both of them at same time.

OpenStudy (anonymous):

where did i mix them?? i have defined Zm,then defined addition and multiplication modulo and have given example 4 each...

OpenStudy (experimentx):

let's work out one at each ... we do with (Zm,+) first.

OpenStudy (anonymous):

ohhh... u mean i have not seperately done it..ryt??? ok got it.... i have jst given it together for my ease....

OpenStudy (experimentx):

but why do you want to give example?

OpenStudy (anonymous):

definition with example is most expected... so that one can easily understand it....

OpenStudy (experimentx):

choose m=10, or 5 for sake of ease. then {0,1,2,3,4} are the elements of group!! show that it remains closed, and is associative and has identity '0', and has inverse.

OpenStudy (anonymous):

then is my answer wrong????

OpenStudy (experimentx):

no i don't think so ... you put stuff in one place, and it hurt my eyes.

OpenStudy (anonymous):

is it ok now??? Let Zm be the set of nonnegative integers less than m: {0,1, ., m−1} 1)The operation +m is defined as a +m b = (a + b) mod m. This is addition modulo m.(Zm,+) Example: Find 7 +11 9 Solution: Using the definitions above: 7 +11 9 = (7 + 9) mod11 = 16 mod11 = 5 2)The operation ∙m is defined as a ∙m b = (a + b) mod m. This is multiplication modulo m. Example: Find 7 ·11 9. Solution: Using the definitions above: 7 ·11 9 = (7 ∙ 9) mod11 = 63 mod 11 = 8

OpenStudy (anonymous):

@experimentX

OpenStudy (experimentx):

sorry ... i was busy!!

OpenStudy (anonymous):

if possible make some alterations and input the needed stuffs... and edit my answer...pls:)

OpenStudy (experimentx):

okay .. you showed that it is closed. you need to test for these http://en.wikipedia.org/wiki/Group_(mathematics)#Definition

OpenStudy (anonymous):

shud i show associativity,identity element and inverse element????

OpenStudy (anonymous):

@experimentX

OpenStudy (experimentx):

yes!! you need to show them!!

OpenStudy (anonymous):

bt in the link u gave... they have shown for multiplication modulo only.... what can i do for addition modulo???????

OpenStudy (experimentx):

same ... show that (a+b) mod m is in the set!!.

OpenStudy (experimentx):

instead of multiplying, we add ,,, since operation is +.

terenzreignz (terenzreignz):

Never seen a Group Theory question before... this should be fun... :D

terenzreignz (terenzreignz):

Once you've shown closure, next step is to show that there is an element e of Zm (a non-negative integer less than m) such that for any k in Zm e + k = k + e = k

OpenStudy (anonymous):

similar to that??? r u sure???

terenzreignz (terenzreignz):

+ here is + modulo m, ok

terenzreignz (terenzreignz):

Quite sure.

terenzreignz (terenzreignz):

Okay, I'll do this (because it's easier :P) and you do the "existence of inverse" part. Let k be in Zm Consider 0. 0 is a non-negative integer. Since m is positive, 0 < m So 0 is in Zm 0 + k = 0 + k (mod m) = k(mod m) = k, since k is in Zm (and is therefore a nonnegative integer less than m). Similarly, k + 0 = k + 0 (mod m) = k(mod m) = k. Thus, we have shown that 0 + k = k + 0 = k Hence, 0 is the identity element in <Zm, +>.

OpenStudy (anonymous):

what abt inverse???

terenzreignz (terenzreignz):

-.- Here's the gist of it. Consider an arbitrary element of Zm, and show that there is something that you can add (modulo m) to it so that the sum is 0 (the additive identity).

OpenStudy (experimentx):

show that for a, m-a is the it's inverse, and m-a is also in the set.

terenzreignz (terenzreignz):

@experimentX A spoiler alert, next time? LOL JK

OpenStudy (experimentx):

well .. sorry!!

OpenStudy (anonymous):

why???

terenzreignz (terenzreignz):

Never mind, lol... Anyway, let a be in Zm. Your task is to show that an element m-a exists, and m-a is in Zm, and a + (m - a) = 0, where "+" here is + (mod m)

OpenStudy (anonymous):

why m-a??? i dnt get it...

terenzreignz (terenzreignz):

Well, remember that in the world of Zm, every multiple of m is basically zero. So, when I said "what do you need to add to a so that you get zero", I really mean what do you need to add to a to get m...

terenzreignz (terenzreignz):

Okay, first of all, the inverse of the additive identity is itself... 0 + 0 = 0, so, yeah... Now, let a be in Zm, where a is not equal to zero. Consider m-a. since a is in Zm, a < m a - a < m - a m - a > 0, meaning m-a is non-negative, (positive, even) m < m + a m - a < m + a - a m - a < m, meaning m-a is less than m. Thus, m-a is non-negative and is less than m, therefore m-a is in Zm. Consider a + (m-a) = a + m - a = m = 0(mod m) Thus, a + (m-a) = 0(mod m) a + (m-a) = 0, in Zm (m-a) is the inverse of a. Thus we have shown that for any element a in Zm, there is an element (m-a) such that a + (m-a) = 0, and therefore, all elements in <Zm, +> have inverses :D

terenzreignz (terenzreignz):

And there you go, we've shown closure, associativity, existence of an identity element and existence of inverses, we can conclude that <Zm, +(mod m)> forms a group. (An Abelian group, too, but more on some other time, I guess)

OpenStudy (anonymous):

great!!!!!!!!!!!!!!

terenzreignz (terenzreignz):

I have to go now, work on the other one: <Zm, *(mod m)> But since we're having spoilers anyway, it's not a group :P The reason for that is for you to find... *cough*oneisthemultiplicativeidentity*cough*theresnothingyoucanmultiplytozerotogetone *cough*zerohasnomultiplicativeinverse *cough* LOL anyway Terence out :D